(no subject)

Emmanuel Monnier (monnier@hep.uchicago.edu)
Fri, 4 Jun 1999 15:43:33 -0500 (CDT)

Shift Summary for day shift, 06/04/99

On Shift: Bob H
Emmanuel

Hours of beam delivered at all intensities: 3.5
Hours of beam delivered at nominal intensity (1.2E12): 3.5
Hours of E799 data: 3.5
Hours of special runs: 0

Accounting of lost time: DA problems, stop run at 10:40 to switch
back to E832 mode.

===================================================================
Runs ended this shift:

13346 - BEAM_799 3.29 M (100 MeV Kick)
13346 - BEAM_799 2.1 M (200 MeV Kick)

Tapes brought to FCC: 10, UP0001 - UP0010 (50% full - E799 data)

===================================================================
Accesses during shift: Many after 10:40

===================================================================
Description of shift:

run time comment
-----------------------
13346 07:00 Inherit run from owl shift, no beam.
" 07:30 getting beam. 1.09 E12.
" 07:45 stream 5 problems, Reset laser repeater (power cycle).
" 08:20 ktevan errors on steam 5, kfrend.
" 08:55 Csi 247 gain drift 7%
" 09:02 Stop run 13346 3.29 M events to tape.
change magnet setting to 800 amps -> 200 MeV Kick.
13347 09:06 Sart run at 200 Mev Kick.
13347 10:40 end runs 2.1 M events.

Called main control to make open access for the next 24 houres.

Upload raw E799 data tapes UP0001 to UP0010 50 % full.

Turned off NM4AN magnet.

Bud and brian are working to free DC2 from helium bag for an access.

No laser scan successfully taken by Peter shanahan due to lots of
ktevana errors.

11:45 Dart people are there to rebuild kernels for all 4 stations.
zillion of problems occured....

3:00 all TRD are out of beam.

3:30 finally everything is back in operation.